WileyPLUS Assignment> Open Assignment BACK Chapter 03, Problem 021 a Chapter 03. Concept An ant, crazed by the Sun on a hot Texas afternoon, darts over an xy plane scratched in the dirt. The x and y components of four consecutive darts are the following, all in centimeters: (27.0, 36.0). (b, -72.0), (-23.0, c)(83.0, -65.0).The overall displacement of the four darts has the xy components (-170, -33.0).What are (a) b, and (b) cy? What are the (c) magnitude and (d) angle (counterclockwise from the positive direction of the x axis) of the overall displacement? Question 02 eater a3. Concept Chapter 93 Concept Question 1 Question 12 003 208 2 Chanter 03 Concent (a) Number (b) Number (c) Number (d) Number Click if you would like to Show Work for this question: Unit Unit l Chapter 03. Problem Unit Open Show Work 022 Review Score Review Results by Study O SAMPLE PROBLEM LENK By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructor Question Attempts: 0 of 6 used SAVE FOR LATERSUBMET ANSWER O Type here to search
Answer #1

answer ) a) here we have

27+bx-23-83=-170

bx=-170+79=91

answer is 91 cm

b) here we have

36-72+cy-65=-33

cy=-33+101=68 cm

so the answer is 68 cm.

c) here we have

R=\sqrt{}-1702+(-33)2=173.17 cm

so the answer is 173 cm or 173.1 cm or 173.17cm.

d) here we have the angle

\Theta=180o+tan-1(-33/-170)=190.990

so the answer is 191o or 190.99o